OpenStudy (anonymous):

The volume of a prism is 35 cm3. What is the volume of this prism, measured in mm3? A. 35,000 mm3 B. 3.5 mm3 C. 0.035 mm3 D. 350 mm3

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

1 cm=10 mm 1 cm^3=10^3 mm^3=1000 mm^3

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 1cm\qquad &10mm\\ 1cm^2&10^2mm^2\\ 1cm^3&10^3mm^3\\ 35cm^3&35\cdot 10^3mm^3 \end{array} \)

OpenStudy (anonymous):

whats the answer